OpenStudy (anonymous):

Use a double-angle identity to find the exact value of each expression.cos4π/3; this is what I got then I was stuck. cos2(2π/3)=1-2sin^2theta but I don't know what sin^2(2π/3) is.

OpenStudy (anonymous):

@ajprincess please help me :)

OpenStudy (ajprincess):

sin(2pi/3)=sqr3/2 so sin^2(2pi/3)=(sqrt3/2)^2=3/4

OpenStudy (anonymous):

ahh thank you!

OpenStudy (ajprincess):

sin(2pi/3)=sin120

OpenStudy (ajprincess):

yw.

OpenStudy (anonymous):

so cos(4pi/3) = cos2(2pi/3)= 1-2sin^2(2pi/3)=1-2(3/4)=-1/2?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

i have another one if you dnt mind?

OpenStudy (ajprincess):

sure post t

OpenStudy (anonymous):

^^ thx; use the half angle identity to find sin(5pi/6); this is how far i got. sin(5pi/6)=sin[(10pi/12)/2] then i was lost. :s

OpenStudy (anonymous):

write an equation in slope intercept form of the line that satisfies the followin condition. x-intercept 12 and y-intercept 14

OpenStudy (ajprincess):

This may be of some use to u.

OpenStudy (ajprincess):

sin(5pi/6)=sin2(5pi/12) see if u can do nw.

OpenStudy (anonymous):

okay, lemme try :)

OpenStudy (anonymous):

@ajprincess how did you get sin(5pi/6)=sin2(5pi/12)? because if I look at the half-angle identity for sin it says sin(x/2)=\[\pm \sqrt({1+cosx})/2\] (5pi/6) x2 = 10pi/6=5pi/3

OpenStudy (ajprincess):

am so sorry. i thought it as double angle identity

OpenStudy (anonymous):

its okay, I found the answer: my sin(5pi/6) x 2 was wrong in the first one :/ silly mistake

OpenStudy (ajprincess):

fine then.